OpenStudy (wintersuntime):

Find the x-intercepts of the following quadratic functions 1.y=x^2+3x-10 2.y=(x-2)^2-9 3.y=1/2(x-1)^2-2 4.y=6x^2+7x-20

OpenStudy (anonymous):

x= -5,2 x= 5, -1 x= 2.73, -0.73 x= 4/3, -5/2

OpenStudy (marcelie):

i think u might have to factor or do quadratic equation

OpenStudy (wintersuntime):

ok thanks

OpenStudy (wintersuntime):

I'll try that

OpenStudy (marcelie):